OpenStudy (bernalcheryl):

How to use Cramer’s rule to solve the following system of equations 5x+3y=7 4x + 5y =3

OpenStudy (phi):

can you make the matrix of coefficients ?

OpenStudy (bernalcheryl):

no

OpenStudy (phi):

write the numbers in front of the x and y to make a row

OpenStudy (sshayer):

\[\left[\begin{matrix}5 & 3 \\ 4 & 5\end{matrix}\right]\left[\begin{matrix}x \\ y \end{matrix}\right]=\left[\begin{matrix}7 \\ 3\end{matrix}\right]\]

OpenStudy (phi):

do you know how to find the determinant of a 2 x 2 matrix ?

OpenStudy (bernalcheryl):

no all this is new for me

OpenStudy (phi):

for a 2 x 2 matix A B C D the determinant is AD - BC can you find the determinant of the matrix ?

OpenStudy (bernalcheryl):

no

OpenStudy (phi):

sshayer wrote down the 2 x 2 matrix do you see it ?

OpenStudy (sshayer):

\[or~AX=B\] \[where ~A=\left[\begin{matrix}5 & 3 \\ 4 & 5\end{matrix}\right]\] \[\left| A \right|=\left|\begin{matrix}5 & 3 \\ 4 & 5\end{matrix}\right|=5*5-4*3=25-12=13\] \[\left| A _{1} \right|=\left|\begin{matrix}7 & 3 \\ 3 & 5\end{matrix}\right|=7*5-3*3=35-9=26\] \[\left| A _{2} \right|=\left[\begin{matrix}5 & 7 \\ 4 & 3\end{matrix}\right]=5*3-4*7=15-28=-13\] \[x=\frac{ \left| A _{1} \right| }{ \left| A \right| }=\frac{ 26 }{ 13 }=2\] \[y=\frac{ \left| A _{2} \right| }{ \left| A \right| }=\frac{ -13 }{ 13 }=-1\]

OpenStudy (bernalcheryl):